# Modular Periodicity of Random Initialized Recurrences

### v3

@@ Abstract
Classical studies of the Fibonacci sequence focus on its periodicity modulo $m$ (the Pisano periods) with canonical initialization. We investigate instead the complete periodic structure arising from all $m^{2}$ possible initializations in $\mathbb{Z}/m\mathbb{Z}$. We discover perfect chiral symmetry between the Fibonacci recurrence $a_{n}=a_{n-1}+a_{n-2}$ and its parity transform $a_{n}=-a_{n-1}+a_{n-2}$ and observe fractal self-similarity in the extension from prime to prime power moduli. Additionally, we classify prime moduli based on their quadratic reciprocity and demonstrate that periodic sequences exhibit weight preservation under modular extension. Furthermore, we define a minima distribution $P(n)$ governed by Lucas ratios, which satisfies the symmetric relation P(n)=P(1n)P(n)=P(1-n). For cyclotomic recurrences, we propose explicit counting functions for the number of distinct periods with connections to necklace enumeration. These findings imply potential connections to Viswanath’s random recurrence, modular forms and L-functions.
Keywords— Fibonacci sequence, Pisano period, cyclotomic polynomial, chiral symmetry, random recurrence, modular arithmetic
@@ 1 Introduction
as shown in [1]. The most celebrated example is the Fibonacci sequence $\{0,1,1,2,3,5,8,13,\ldots\}$, generated by $x^{2}-x-1$ with initial values $a_{0}=0,a_{1}=1$. This sequence grows exponentially, but when reduced modulo $m$, it becomes periodic with period length $\pi(m)$, known as the Pisano period [2]. Determining $\pi(m)$ explicitly remains an open problem in number theory.
@@
To approach this systematically, we focus on recurrences with restricted coefficients, namely monic quadratic polynomials with coefficients in $\{-1,+1\}$ (also known as Littlewood polynomials [3]). Two of these four polynomials are $x^{2}-x-1$ and $x^{2}+x-1$, which generate the Fibonacci recurrence $a_{n}=a_{n-1}+a_{n-2}$ and its parity transform $a_{n}=-a_{n-1}+a_{n-2}$. The other two are the cyclotomic polynomials $\Phi_{3}(x)=x^{2}+x+1$ and $\Phi_{6}(x)=x^{2}-x+1$, whose roots are primitive roots of unity. Interestingly, the Fibonacci recurrence (which is the negation from $\Phi_{3}$) and the recurrence $a_{n}=a_{n-1}-a_{n-2}$ (from $\Phi_{6}$) form the two deterministic branches of Viswanath’s random recurrence $a_{n}=a_{n-1}\pm a_{n-2}$, which exhibits exponential growth with rate approximately $1.13198824$ [4]. This connection between non-cyclotomic and cyclotomic recurrences motivates our parallel investigation of both families. More generally, the $n$th cyclotomic polynomial is defined as
@@
satisfying $\prod_{d\mid n}\Phi_{d}(x)=x^{n}-1$. These polynomials are fundamental in algebra and number theory, yet their role in modular periodicity of recurrences remains largely unexplored. Since their roots lie on the unit circle, cyclotomic recurrences exhibit fundamentally different behavior than exponentially growing sequences; they remain bounded, making them a natural starting point for classification.
This paper investigates the period landscapes of both cyclotomic and non-cyclotomic recurrences across three main sections. In Section 2, we conjecture explicit counting functions for the number of distinct periods modulo $m$ for various cyclotomic families, including $\Phi_{p}$, $\Phi_{2p}$, $\Phi_{p^{j}}$ and $x^{n}-1$, all supported by extensive computational verification. In Section 3, the non-cyclotomic Fibonacci and parity recurrences are explored, revealing unexpected chiral symmetry. Despite having algebraically distinct characteristic roots, these recurrences produce identical period structures for all tested moduli. We classify their prime moduli via the Legendre symbol and a parameter $\alpha$ governing both period counts and root orders in field extensions, observe fractal self-similarity at prime power moduli and conjecture weight preservation under modular extension. Furthermore, we establish that both recurrences share an identical minimum distribution governed by Lucas number ratios. In Section 4, we discuss broader implications and open questions and suggest that deterministic periodicity in finite rings has potential relations to stochastic growth in random recurrences and to classical number-theoretical objects such as modular forms and L-functions.
@@ 2 Periodicity of Cyclotomic Recurrences
is the number of different aperiodic $m$-ary necklaces of length $r$, as presented in [5].It is also the number of monic irreducible polynomials of degree r over a finite field $\mathbb{F}_{q}$ (see [6]), with $\mu$ being the classic Möbius function. $M(m,r)$ also refers to Moureau’s necklace-counting function or MacMahon’s formula.
@@ 3 Periodicity of Fibonacci Recurrences
We now turn our focus to the Fibonacci recurrence $a_{n}=a_{n-1}+a_{n-2}$ and its parity transform $a_{n}=-a_{n-1}+a_{n-2}$ with random integer initialization $(a_{0},a_{1})\in(\mathbb{Z}/m\mathbb{Z})^{2}$. The implied mirror-symmetry can be seen in Table 1 for all moduli $m$, where the number of periods and lengths are similar for both recurrences, and where every period corresponds to a chiral period from the other recurrence.
@@
Consider the Fibonacci recurrence $a_{n}=a_{n-1}+a_{n-2}$ and its chiral recurrence $a_{n}=-a_{n-1}+a_{n-2}$ with random integer initialization $\{a_{0},a_{1}\}\in\mathbb{Z}^{2}$. Although these sequences diverge as $n\to\pm\infty$, they possess well-defined absolute minima. Let $F_{n}$ and $L_{n}$ denote the $n$-th Fibonacci and Lucas numbers respectively. Define $P(n)$ as the probability that a randomly initialized sequence has its absolute minimum at position $n$. Then:
@@ 4 Perspectives
This paper examined the modular periodicity of randomly initialized Fibonacci and cyclotomic recurrences, revealing previously unknown symmetries. We conjectured explicit counting formulas for cyclotomic recurrences and discovered perfect chiral symmetry between the Fibonacci and parity recurrences. We classified prime moduli based on the Legendre symbol and a parameter $\alpha$ governing both period counts and multiplicative orders of roots in field extensions. Additionally, we observed fractal self-similarity at prime power moduli, established weight preservation under modular extension, and derived a probability distribution for sequence minima governed by Lucas ratios.
The framework of random initialization naturally extends to higher-degree polynomials. For example, the two order-6 recurrences $a_{n}=\pm a_{n-3}+a_{n-6}$ (negations from $\Phi_{9}$ and $\Phi_{18}$) exhibit chiral symmetry analogous to the Fibonacci case for all tested moduli $m\leq 10$, though their period structure is more complex with up to five distinct lengths for prime moduli. Whether the classification principles established for quadratic recurrences extend to higher-order systems remains an open question.
Preliminary computation reveals systematic embedding of Fibonacci periods into cyclotomic recurrences: class A periods appear in $\Phi_{2(p+1)p^{j}/\alpha}$ modulo $p$, while class B1 periods appear in $\Phi_{(p-1)p^{j}/\alpha}$ modulo $p$. Class B2 primes exhibit more complex behavior, with some showing additional projections to composite moduli. These patterns suggest $\alpha$ encodes structural information about cyclotomic decompositions beyond what the Legendre symbol reveals. Investigating embeddings into other polynomial families — such as Fekete polynomials, which relate to cyclotomic polynomials and L-functions [7] — could reveal universal principles governing recurrence periodicity in finite rings.

### v4

@@ Abstract
Classical studies of the Fibonacci sequence focus on its periodicity modulo $m$ (the Pisano periods) with canonical initialization. We investigate instead the complete periodic structure arising from all $m^{2}$ possible initializations in $\mathbb{Z}/m\mathbb{Z}$. We discover perfect mirror symmetry between the Fibonacci recurrence $a_{n}=a_{n-1}+a_{n-2}$ and its parity transform $a_{n}=-a_{n-1}+a_{n-2}$ and observe fractal self-similarity in the extension from prime to prime power moduli. Additionally, we classify prime moduli based on their quadratic reciprocity and demonstrate that periodic sequences exhibit weight preservation under modular extension. Furthermore, we define a minima distribution $P(n)$ governed by Lucas ratios, which satisfies the symmetric relation P(n)=P(1n)P(n)=P(1-n). For cyclotomic recurrences, we propose explicit counting functions for the number of distinct periods with connections to necklace enumeration. These findings imply potential connections to Viswanath’s random recurrence, modular forms and L-functions.
Keywords— Fibonacci sequence, Pisano period, cyclotomic polynomial, random recurrence, modular arithmetic, mirror symmetry
@@ 1 Introduction
as shown in [undef]. The most celebrated example is the Fibonacci sequence $\{0,1,1,2,3,5,8,13,\ldots\}$, generated by $x^{2}-x-1$ with initial values $a_{0}=0,a_{1}=1$. This sequence grows exponentially, but when reduced modulo $m$, it becomes periodic with period length $\pi(m)$, known as the Pisano period [undefa]. Determining $\pi(m)$ explicitly remains an open problem in number theory.
@@
To approach this systematically, we investigate two families of recurrences that exhibit complementary structure. The first family consists of cyclotomic polynomials $\Phi_{n}(x)$, defined as
@@
satisfying $\prod_{d\mid n}\Phi_{d}(x)=x^{n}-1$. These polynomials are fundamental in algebra and number theory, yet their role in modular periodicity remains largely unexplored. Since all their roots are simple and lie on the unit circle, cyclotomic recurrences remain bounded, making them a natural starting point for classification.
The second family consists of exponentially growing sequences, exemplified by the Fibonacci recurrence $a_{n}=a_{n-1}+a_{n-2}$ (from $x^{2}-x-1$) and its parity transform $a_{n}=-a_{n-1}+a_{n-2}$ (from $x^{2}+x-1$). Interestingly, these two form the deterministic branches of Viswanath’s random recurrence $a_{n}=\pm a_{n-1}+a_{n-2}$, which also exhibits exponential growth with rate approximately $1.13198824$ [undefb]. Moreover, they arise as sign-reversed versions of the cyclotomic recurrences from $\Phi_{3}(x)=x^{2}+x+1$ and $\Phi_{6}(x)=x^{2}-x+1$. This connection between unit-root-bounded and exponentially growing recurrences through coefficient sign reversal motivates our parallel investigation of both families.
This paper investigates the period landscapes of both cyclotomic and non-cyclotomic recurrences across three main sections. In Section 2, we conjecture explicit counting functions for the number of distinct periods modulo $m$ for various cyclotomic families, including $\Phi_{p}$, $\Phi_{2p}$, $\Phi_{p^{j}}$ and $x^{n}-1$, all supported by extensive computational verification. In Section 3, the non-cyclotomic Fibonacci and parity recurrences are explored, revealing unexpected mirror symmetry. Despite having algebraically distinct characteristic roots, these recurrences produce identical period structures for all tested moduli. We classify their prime moduli via the Legendre symbol and a parameter $\alpha$ governing both period counts and root orders in field extensions, observe fractal self-similarity at prime power moduli and conjecture weight preservation under modular extension. Furthermore, we establish that both recurrences share an identical minimum distribution governed by Lucas number ratios. In Section 4, we discuss broader implications and open questions and suggest that deterministic periodicity in finite rings has potential relations to stochastic growth in random recurrences and to classical number-theoretical objects such as modular forms and L-functions.
@@ 2 Periodicity of Cyclotomic Recurrences
is the number of different aperiodic $m$-ary necklaces of length $r$, as presented in [undefc].It is also the number of monic irreducible polynomials of degree r over a finite field $\mathbb{F}_{q}$ (see [undefd]), with $\mu$ being the classic Möbius function. $M(m,r)$ also refers to Moureau’s necklace-counting function or MacMahon’s formula.
@@ 3 Periodicity of Fibonacci Recurrences
We now turn our focus to the Fibonacci recurrence $a_{n}=a_{n-1}+a_{n-2}$ and its parity transform $a_{n}=-a_{n-1}+a_{n-2}$ with random integer initialization $(a_{0},a_{1})\in(\mathbb{Z}/m\mathbb{Z})^{2}$. The implied mirror symmetry can be seen in Table 1 for all moduli m, where the number of periods and lengths are similar for both recurrences, and where every period corresponds to a mirror period from the other recurrence.
@@
Consider the Fibonacci recurrence $a_{n}=a_{n-1}+a_{n-2}$ and its parity recurrence $a_{n}=-a_{n-1}+a_{n-2}$ with random integer initialization $\{a_{0},a_{1}\}\in\mathbb{Z}^{2}$. Although these sequences diverge as $n\to\pm\infty$, they possess well-defined absolute minima. Let $F_{n}$ and $L_{n}$ denote the $n$-th Fibonacci and Lucas numbers respectively. Define $P(n)$ as the probability that a randomly initialized sequence has its absolute minimum at position $n$. Then:
@@ 4 Perspectives
This paper examined the modular periodicity of randomly initialized Fibonacci and cyclotomic recurrences, revealing previously unknown symmetries. We conjectured explicit counting formulas for cyclotomic recurrences and discovered perfect mirror symmetry between the Fibonacci and parity recurrences. We classified prime moduli based on the Legendre symbol and a parameter $\alpha$ governing both period counts and multiplicative orders of roots in field extensions. Additionally, we observed fractal self-similarity at prime power moduli, established weight preservation under modular extension, and derived a probability distribution for sequence minima governed by Lucas ratios.
The framework of random initialization naturally extends to higher-degree polynomials. For example, the two order-6 recurrences $a_{n}=\pm a_{n-3}+a_{n-6}$ (negations from $\Phi_{9}$ and $\Phi_{18}$) exhibit mirror symmetry analogous to the Fibonacci case for all tested moduli $m\leq 11$, though their period structure is more complex with up to five distinct lengths for prime moduli. Whether the classification principles established for quadratic recurrences extend to higher-order systems remains an open question.
Preliminary computation reveals systematic embedding of Fibonacci periods into cyclotomic recurrences: class A periods appear in $\Phi_{2(p+1)p^{j}/\alpha}$ modulo $p$, while class B1 periods appear in $\Phi_{(p-1)p^{j}/\alpha}$ modulo $p$. Class B2 primes exhibit more complex behavior, with some showing additional projections to composite moduli. These patterns suggest $\alpha$ encodes structural information about cyclotomic decompositions beyond what the Legendre symbol reveals. Investigating embeddings into other polynomial families — such as Fekete polynomials, which relate to cyclotomic polynomials and L-functions [undefe] — could reveal universal principles governing recurrence periodicity in finite rings.